OpenStudy (anonymous):

The semicircles are congruent to one another with a diameter of 4 inches. What is the perimeter of the sign?

OpenStudy (unklerhaukus):

the perimeter is equal to the sums of the lengths of the two straight lines plus the circumference of three three circles (from 6 semicircles)

OpenStudy (unklerhaukus):

the circumference of a circle is \[2\pi r\]

OpenStudy (anonymous):

so how would i plug the numbers in ?

OpenStudy (unklerhaukus):

\[2\times20 \text{in}+6\times2\pi\times\left(\frac {4\text {in} }2\right)\]

OpenStudy (anonymous):

i got 64 i no its wrong

OpenStudy (anonymous):

do i multiply 3.14?

OpenStudy (unklerhaukus):

if you need a numerical answer you can use \[\pi\approx 3.142\]

OpenStudy (anonymous):

now i got 144 wat am i doing wrong?

OpenStudy (unklerhaukus):

\[=40\text {in}+24\pi\text{ in}\] \[=40+(24\times3.14)\text {in}\]

OpenStudy (anonymous):

ooo i gt 38 ?

OpenStudy (unklerhaukus):

how could you et less than 40?

OpenStudy (unklerhaukus):

*get

OpenStudy (anonymous):

from the first problem u pt up i used 3.14*2

OpenStudy (unklerhaukus):

huh?

OpenStudy (anonymous):

the third post u put up

OpenStudy (anonymous):

i followed the equation. i think its rite because these are the options : about 38 inches about 78 inches about 80 inches about 76 inches

OpenStudy (unklerhaukus):

\[P=2l+3\times2\pi r=2l+3\times2\pi \frac d2\]

OpenStudy (unklerhaukus):

sorry i got confused withe the semicircles before